Sol Summary – April 9th

Crew 192 Sol 1 Summary Report

Sol 1

Summary Title: PHEnOM Blue Crew Has Landed

Author: Victoria Varone

Mission Status: Ground Control to Major Tom, PHEnOM Blue Crew 192 has landed at the Mars Desert Research Station. Unfortunately we’ve “lost” a crew-member during reentry to Mars. Our science adviser was accidentally jettisoned along with many of our scientific experiments. We have renamed our habitat the Bowie Base in honor of the inspiration from the “star Man” himself David Bowie. Our team’s mission is to overcome this adversity and persevere in the name of science. We have good intelligence that some of the projects may have survived reentry so we will be performing EVAs to scout possible locations where the science may have re-entered. We intend to perform pilot studies and other scientific information gathering techniques as well.

Sol Activity Summary: We started our day by making sure we have an adequate water supply for our mission. Following that, we scouted the area around our habitat and took some photos for documentation. We also checked on our remaining science experiments like the locking mechanism of our observatory, which eventually opened after several unsuccessful attempts.

Next, we tested our ATVs and rovers to make sure they’re working properly for our EVAs and scouting missions. It was interesting getting to use them on the dusty, sandy Martian landscape.

After our test, we headed back inside the hab for our first crew dinner during our mission, and after that, two crew members conducted our first short-range EVA to test our spacesuits and communications devices. Upon their return, our GreenHab Officer went out to water the plants and our XO conducted a short engineering EVA around the hab for some repairs.

Look Ahead Plan: As we continue to search for our lost science experiments, we’ll stretch our EVAs further and further out into the Martian landscape each day. We’re also experimenting with various types of dehydrated “space food” to see what brings us the best combination of nutrition, convenience, and satisfaction. Some of the dehydrated items have proven to be more difficult to cook with than expected.

Anomalies in Work: None

Weather: The weather was beautiful all day today, clear skies with very few clouds. We’re getting a good feel for the desert climate, it’s dryness, and its variation in temperature. It seems to be somewhat chilly in the morning, warmer by midday (even a bit hot), and cools down nicely at night, sometimes even becoming very cold during the night itself. The dry air is also a change for some of the crew.

Crew Physical Status: All crew are physically healthy and safe.

EVA: Today two crewmembers conducted one short-range EVA that lasted approximately one hour. See additional EVA report.

Reports to be filed: Sol Summary Report, HSO Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: We would very much appreciate it if mission control could send us forecast sunset times (actual, civil, and nautical) for each day of our mission, which will allow us some better planning for EVAs.

Copyright © The Mars Society. All rights reserved. | Main Site